# Recursive decoding of projective Reed-Muller codes

### v1

@@ 3.1. Projective Reed-Solomon codes
Therefore, $\operatorname{wt}(e_{2})=0$ and $c_{2}=r_{2}$. Let $f\in\mathbb{F}_{q}[x_{0},x_{1}]_{d}$ such that $\operatorname{ev}^{\mathbb{P}}_{{1}}(f)=c$. From this we obtain that the coefficient of $x_{1}^{d}$ in $f$ is $r_{2}$. Moreover, $g=(f-r_{2}x_{1}^{d})(1,x_{1})\in\mathbb{F}_{q}[x_{1}]_{\leq d-1}$. Thus,
@@ 3.2. Projective Reed-Muller codes over the projective plane
In this subsection we generalize the previous decoding algorithm to $\operatorname{PRM}_{d}(2)$. Let $m=2$, $\nu(q-1)<d\leq(\nu+1)(q-1)$, $0\leq\nu\leq 1$, and consider
@@
$u=\operatorname{ev}^{\mathbb{A}}_{{2}}(f^{0}_{\leq d-1}),\;v=\operatorname{ev}^{\mathbb{P}}_{{1}}(f^{0}_{d}),\;c=\operatorname{ev}^{\mathbb{P}}_{{2}}(h^{d}_{0}(f^{0})).$
@@
When decoding, we do not know if $\operatorname{wt}(e_{1})<t_{1}$, but we can nonetheless proceed as above and obtain a vector $c^{\prime}$. If the corresponding error vector $(r-c^{\prime})$ has weight less than $t$, then we conclude $c=c^{\prime}$ and we finish the algorithm. If not, then we know $\operatorname{wt}(e_{1})\geq t_{1}$. In that case, we have $\operatorname{wt}(e_{2})<t_{2}$ and $\operatorname{wt}(e_{2})$ is within the error correction capability of $\operatorname{PRS}_{d}$, which means that we can consider $(v,g)=D^{\mathbb{P}}_{d}(1)(v+e_{2})$ (using the algorithm from Subsection 3.1). If $d=q-1$, $t_{2}=1$, $\operatorname{wt}(e_{2})=0$, we assume that $D^{\mathbb{P}}_{d}(1)$ just returns the received vector and its corresponding polynomial. From $v$, we obtain $v_{\xi,d}$, and we can consider
@@
In order to obtain $f$, we need to use the information from the last $q+1$ coordinates. First we decompose $f^{0}$ as follows (recall Definition 2.7):
@@ 3.3. Projective Reed-Muller codes, general case
Let $d-1=\nu(q-1)+\mu$, $d\leq m(q-1)$, $10\leq\mu<q-$ 1. Let $r=c+e$ be a received codeword, with $c\in\operatorname{PRM}_{d}(m)$ and $\operatorname{wt}(e)<t_{0}$. Then Algorithm 1 can be used to recover $c$ from $r$.
@@
$u=\operatorname{ev}^{\mathbb{A}}_{{m}}(f^{0}_{\leq d-1}),\;v=\operatorname{ev}^{\mathbb{P}}_{{m-1}}(f^{0}_{d}),\;c=\operatorname{ev}^{\mathbb{P}}_{{m}}(h^{d}_{0}(f^{0})).$
Therefore, we have obtained $(c,f)=((u+v_{\xi,d},v),h^{d}_{0}(f^{0}))$.
@@
Since when decoding we do not know if $\operatorname{wt}(e_{1})<t_{1}$, in the first part of Algorithm 1 we follow the previous procedure to obtain a vector $c^{\prime}$. Then we check if $\operatorname{wt}(r-c^{\prime})<t$. If that is the case, we return $c^{\prime}$ and the corresponding polynomial $f^{\prime}$. If not, we proceed to the second part of Algorithm 1.
Second part: what follows corresponds to the second part of Algorithm 2, which starts at Line 31. Since the first part of Algorith 1 has failed, it means that $\operatorname{wt}(e_{1})\geq t_{1}$, and therefore $\operatorname{wt}(e_{2})<t_{2}=\eta_{d}(m-1)/2$. By the induction hypothesis, we can consider $(v,g)=D^{\mathbb{P}}_{d}(m-1)(v+e_{2})$, and then perform the decoding (recall Remark 2.5)
@@
31:$(v,g)\leftarrow D^{\mathbb{P}}_{d}(m-1)(r_{2})$triangleright Second part
32:$v_{\xi,d}=(v,\xi^{d}v,\xi^{2d}v,\dots,\xi^{(q-2)d}v,0)$
33:$(u,f^{0}_{\leq d-1})\leftarrow D^{\mathbb{A}}_{d-1}(m)(r_{1}-v_{\xi,d})$
34:$f=h_{d}^{0}(f^{0}_{\leq d-1})+g$
35:$c=(u+v_{\xi,d},v)$
36:return $(c,f)$
4. Analysis of the algorithm and improvements
Let $1\leq d\leq m(q-1)$ with $d-1=\nu(q-1)+\mu$, $10\leq\mu<q-$ 1. Algorithm 1 decodes any number of errors lower than $\eta_{d}(m)/2$. The previous algorithm known for PRM codes was given in [26], and it was able to correct any number of errors lower than $\operatorname{wt}(\operatorname{RM}_{d}(m))/2$ (see [26, Cor. 5.2]). Looking at Equation (3), it is clear that
and therefore $\operatorname{wt}(\operatorname{RM}_{d}(m))/2<\eta_{d}(m)/2$, which means that Algorithm 1 can decode more errors in general. Moreover, since $\operatorname{wt}(\operatorname{RM}_{d}(m))<\operatorname{wt}(\operatorname{RM}_{d-1}(m))=\operatorname{wt}(\operatorname{PRM}_{d}(m))$, the algorithm from [26] decodes up to the error correction capability of $\operatorname{PRM}_{d}(m)$ only if $\operatorname{wt}(\operatorname{RM}_{d}(m))=\operatorname{wt}(\operatorname{RM}_{d-1}(m))-1$ (which only happens if and only if $d>(m-1)(q-1)$, i.e., $\nu=m-1$) and $\operatorname{wt}(\operatorname{RM}_{d-1}(m))=\operatorname{wt}(\operatorname{PRM}_{d}(m))$ is even. On the other hand, Algorithm 1 decodes up to the error correction capability of $\operatorname{PRM}_{d}(m)$ whenever $\mu=0$ or $\nu=m-1$ (see Lemma 2.9), independent of the parity of $\operatorname{wt}(\operatorname{PRM}_{d}(m))$, and in some additional cases depending on the parity of $\operatorname{wt}(\operatorname{PRM}_{d}(m))$. In Figures 1 and 2 we show the value of $T_{0}/T$, where
i.e., $T_{0}/T$ is the quotient of the number of errors that Algorithm 1 can correct with respect to the error correction capability of $\operatorname{PRM}_{d}(m)$. For example, for $q=3$ and $m=2$, Algorithm 1 can correct up to the error correction capability of $\operatorname{PRM}_{d}(2)$ for any degree for which $\operatorname{PRM}_{d}(2)$ is non trivial. As noted before, for $\mu=0$ or $\nu=m-1$, we always have $T_{0}/T=1$, i.e., we correct up to the error correction capability of $\operatorname{PRM}_{d}(m)$.
Figure 1. $T_{0}/T$ as a function of $1\leq d\leq m(q-1)$, for $m=2$.
@@ 4. Analysis of the algorithm and improvements
Let $d-1=\nu(q-1)+\mu$, $d\leq m(q-1)$, $10\leq\mu<q-$ 1, and let $e$ be an error vector with $\operatorname{wt}(e)<t$. Assume there is some $i\leq m$ such that
@@
Additionally, some decoders for RM codes are also known to decode more errors than the error correction capability, as long as this errors are in general position (see [6]). By using such decoders with Algorithm 2, again we may also be able to correct more errors than $T_{0}$.
The subfield subcodes of PRM codes (that is, the intersection of a PRM code with ${\mathbb{F}}_{q^{\prime}}^{n}$, where ${\mathbb{F}}_{q^{\prime}}\subset\mathbb{F}_{q}$) have been studied in [14, 32]. They have been shown to have good parameters, and they can be used for practical applications [13, 28, 27]. Since these codes are subcodes of PRM codes, they can also be decoded with Algorithms 1 and 2. The usual problem with this approach is that the subfield subcode is defined over a smaller field ${\mathbb{F}}_{q^{\prime}}$, and the decoding might require operations over the bigger field $\mathbb{F}_{q}$. However, in [32, Cor. 3], it is shown that, if $d$ is a multiple of $(q-1)/(q^{\prime}-1)$, then the recursive construction from Theorem 2.8 also works for the subfield subcodes. Thus, Algorithm 1 can decode the subfield subcodes for these degrees using only operations over ${\mathbb{F}}_{q^{\prime}}$, if we consider appropriate decoders for the subfield subcodes of affine RM codes.
@@ 5. Complexity
This means that the length of $\operatorname{PRM}_{d}(m)$ is about $q/(q-1)$ times higher than that of $\operatorname{RM}_{d}(m)$. This does not change the order of complexity, and therefore we can use both $p_{m}$ or $n^{\mathbb{A}}$ to express the order of complexity of $D^{\mathbb{P}}_{d}(m)$.
@@
As we explained in Section 4, the decoding algorithm from [26] decodes any number of errors lower than $\operatorname{wt}(\operatorname{RM}_{d}(m))/2$. In that setting, Algorithm 1 always finishes in the first part and only requires to use $D^{\mathbb{A}}_{d}(m)$. In this sense, Algorithm 1 generalizes the algorithm from [26] not only because it can correct more errors, but also because their complexity is the same when both algorithms can be applied.

### v2

@@ 3.1. Projective Reed-Solomon codes
Therefore, $\operatorname{wt}(e_{2})=0$ and $c_{2}=r_{2}$. Let $f\in\mathbb{F}_{q}[x_{0},x_{1}]_{d}$ such that $\operatorname{ev}^{\mathbb{P}}_{{1}}(f)=c$. From this, we obtain that the coefficient of $x_{1}^{d}$ in $f$ is $r_{2}$. Moreover, $g=(f-r_{2}x_{1}^{d})(1,x_{1})\in\mathbb{F}_{q}[x_{1}]_{\leq d-1}$. Thus,
@@ 3.2. Projective Reed-Muller codes over the projective plane
In this subsection, we generalize the previous decoding algorithm to $\operatorname{PRM}_{d}(2)$. Let $m=2$, $\nu(q-1)<d\leq(\nu+1)(q-1)$, $0\leq\nu\leq 1$, and consider
@@
$u=\operatorname{ev}^{\mathbb{A}}_{{2}}(f^{0}_{\leq d-1}),\;v=\operatorname{ev}^{\mathbb{P}}_{{1}}(f^{0}_{d}),\;c=\operatorname{ev}^{\mathbb{P}}_{{2}}(h^{0}_{d}(f^{0})).$
@@
When decoding, we do not know if $\operatorname{wt}(e_{1})<t_{1}$, but we can nonetheless proceed as above and obtain a vector $c^{\prime}$. If the corresponding error vector $(r-c^{\prime})$ has weight less than $t$, then we conclude $c=c^{\prime}$ and we finish the algorithm. If not, then we know $\operatorname{wt}(e_{1})\geq t_{1}$. In that case, we have $\operatorname{wt}(e_{2})<t_{2}$ and $\operatorname{wt}(e_{2})$ is within the error correction capability of $\operatorname{PRS}_{d}$, which means that we can consider $(v,g)=D^{\mathbb{P}}_{d}(1)(v+e_{2})$ (using the algorithm from Subsection 3.1). If d=q1d=q-1, $t_{2}=1$, $\operatorname{wt}(e_{2})=0$, we assume that $D^{\mathbb{P}}_{d}(1)$ just returns the received vector and its corresponding polynomial. From $v$, we obtain $v_{\xi,d}$, and we can consider
@@
To obtain $f$, we need to use the information from the last $q+1$ coordinates. First we decompose $f^{0}$ as follows (recall Definition 2.7):
@@ 3.3. Projective Reed-Muller codes, general case
Let $d-1=\nu(q-1)+\mu$, $d\leq m(q-1)$, $0\leq\mu<q-1$. Let $r=c+e$ be a received codeword, with $c\in\operatorname{PRM}_{d}(m)$ and $\operatorname{wt}(e)<t_{0}$. Then Algorithm 1 can be used to recover $c$ from $r$.
@@
$u=\operatorname{ev}^{\mathbb{A}}_{{m}}(f^{0}_{\leq d-1}),\;v=\operatorname{ev}^{\mathbb{P}}_{{m-1}}(f^{0}_{d}),\;c=\operatorname{ev}^{\mathbb{P}}_{{m}}(h^{0}_{d}(f^{0})).$
Therefore, we have obtained $(c,f)=((u+v_{\xi,d},v),h^{0}_{d}(f^{0}))$.
@@
Since we do not know if $\operatorname{wt}(e_{1})<t_{1}$ when decoding, in the first part of Algorithm 1 we follow the previous procedure to obtain a vector $c^{\prime}$. Then we check if $\operatorname{wt}(r-c^{\prime})<t$. If that is the case, we return $c^{\prime}$ and the corresponding polynomial $f^{\prime}$. If not, we proceed to the second part of Algorithm 1.
Second part: what follows corresponds to the second part of Algorithm 2, which starts at Line 31. Since the first part of Algorithm 1 has failed, it means that $\operatorname{wt}(e_{1})\geq t_{1}$, and therefore $\operatorname{wt}(e_{2})<t_{2}=\eta_{d}(m-1)/2$. By the induction hypothesis, we can consider $(v,g)=D^{\mathbb{P}}_{d}(m-1)(v+e_{2})$, and then perform the decoding (recall Remark 2.5)
@@
[…]
36:return $(c,f)$
4. Analysis of the algorithm and improvements
Let $1\leq d\leq m(q-1)$ with $d-1=\nu(q-1)+\mu$, $0\leq\mu<q-1$. Algorithm 1 decodes any number of errors lower than $\eta_{d}(m)/2$. The previous algorithm known for PRM codes was given in [28], and it was able to correct any number of errors lower than $\operatorname{wt}(\operatorname{RM}_{d}(m))/2$ (see [28, Cor. 5.2]). Looking at Equation (3), it is clear that
and therefore $\operatorname{wt}(\operatorname{RM}_{d}(m))/2<\eta_{d}(m)/2$, which means that Algorithm 1 can decode more errors in general. Moreover, since $\operatorname{wt}(\operatorname{RM}_{d}(m))<\operatorname{wt}(\operatorname{RM}_{d-1}(m))=\operatorname{wt}(\operatorname{PRM}_{d}(m))$, the algorithm from [28] decodes up to the error correction capability of $\operatorname{PRM}_{d}(m)$ only if $\operatorname{wt}(\operatorname{RM}_{d}(m))=\operatorname{wt}(\operatorname{RM}_{d-1}(m))-1$ (which only happens if and only if d>(m1)(q1)d>(m-1)(q-1), i.e., $\nu=m-1$) and $\operatorname{wt}(\operatorname{RM}_{d-1}(m))=\operatorname{wt}(\operatorname{PRM}_{d}(m))$ is even. On the other hand, Algorithm 1 decodes up to the error correction capability of $\operatorname{PRM}_{d}(m)$ whenever $\mu=0$ or $\nu=m-1$ (see Lemma 2.9), independent of the parity of $\operatorname{wt}(\operatorname{PRM}_{d}(m))$, and in some additional cases depending on the parity of $\operatorname{wt}(\operatorname{PRM}_{d}(m))$. In Figures 1, 2 and 3 we show the value of $T_{0}/T$, where
i.e., $T_{0}/T$ is the quotient of the number of errors that Algorithm 1 can correct with respect to the error correction capability of $\operatorname{PRM}_{d}(m)$. For example, for $q=3$ and $m=2$, Algorithm 1 can correct up to the error correction capability of $\operatorname{PRM}_{d}(2)$ for any degree for which $\operatorname{PRM}_{d}(2)$ is non trivial. As noted before, for $\mu=0$ or $\nu=m-1$, we always have $T_{0}/T=1$, i.e., we correct up to the error correction capability of $\operatorname{PRM}_{d}(m)$. Also note that, for $q$ large, the error correction capability of the decoding algorithm is very close to that of $\operatorname{PRM}_{d}(m)$ for most degrees (see Figure 2).
For some degrees, PRM codes can be seen as Grassmann codes, for which some decoding algorithms have been studied [6, 23]. However, Algorithm 1 can correct more errors in general than these decoders for the corresponding PRM codes. For example, in [6, Rem. 4.7], the authors mention that their algorithm decodes up to half the minimum distance for $q=2$, and roughly up to $\operatorname{wt}(\operatorname{PRM}_{d}(1))/4$ for large $q$. Algorithm 1 decodes always up to the error correction capability of $\operatorname{PRM}_{d}(1)$ for any $q$.
Figure 1. $T_{0}/T$ as a function of $1\leq d\leq m(q-1)$, for $m=2$ and $q$ small.
${1}$${2}$${3}$${4}$${0.6}$${0.8}$${1.0}$$q=3,m=2$
${1}$${2}$${3}$${4}$${5}$${6}$${0.6}$${0.8}$${1.0}$$q=4,m=2$
${1}$${2}$${3}$${4}$${5}$${6}$${7}$${8}$${0.6}$${0.8}$${1.0}$$q=5,m=2$
${1}$${2}$${3}$${4}$${5}$${6}$${7}$${8}$${9}$${10}$${11}$${12}$${0.6}$${0.8}$${1.0}$$q=7,m=2$
Figure 2. $T_{0}/T$ as a function of $1\leq d\leq m(q-1)$, for $m=2$ and $q$ large.
$1$$5$$10$$15$$20$$25$$30$$0.6$$0.8$$1.0$$q=16,\;m=2$
$1$$20$$40$$60$$80$$100$$120$$0.6$$0.8$$1.0$$q=64,\;m=2$
$1$$100$$200$$300$$400$$500$$0.6$$0.8$$1.0$$q=256,\;m=2$
$1$$300$$600$$900$$1200$$1500$$1800$$0.6$$0.8$$1.0$$q=1024,\;m=2$
@@ 4. Analysis of the algorithm and improvements
Let $d-1=\nu(q-1)+\mu$, $d\leq m(q-1)$, $0\leq\mu<q-1$, and let $e$ be an error vector with $\operatorname{wt}(e)<t$. Assume there is some $i\leq m$ such that
@@
Additionally, some decoders for RM codes are also known to decode more errors than the error correction capability, as long as these errors are in general position (see [7]). By using such decoders with Algorithm 2, again we may also be able to correct more errors than $T_{0}$.
The subfield subcodes of PRM codes (that is, the intersection of a PRM code with ${\mathbb{F}}_{q^{\prime}}^{n}$, where ${\mathbb{F}}_{q^{\prime}}\subset\mathbb{F}_{q}$) have been studied in [15, 34]. They have been shown to have good parameters, and they can be used for practical applications [14, 30, 29]. Since these codes are subcodes of PRM codes, they can also be decoded with Algorithms 1 and 2. The usual problem with this approach is that the subfield subcode is defined over a smaller field ${\mathbb{F}}_{q^{\prime}}$, and the decoding might require operations over the bigger field $\mathbb{F}_{q}$. However, in [34, Cor. 3], it is shown that, if $d$ is a multiple of $(q-1)/(q^{\prime}-1)$, then the recursive construction from Theorem 2.8 also works for the subfield subcodes. Thus, Algorithm 1 can decode the subfield subcodes for these degrees using only operations over ${\mathbb{F}}_{q^{\prime}}$, if we consider appropriate decoders for the subfield subcodes of affine RM codes.
@@ 5. Complexity
This means that the length of $\operatorname{PRM}_{d}(m)$ is about q/(q1)q/(q-1) times higher than that of $\operatorname{RM}_{d}(m)$. This does not change the order of complexity, and therefore we can use both $p_{m}$ or $n^{\mathbb{A}}$ to express the order of complexity of $D^{\mathbb{P}}_{d}(m)$.
@@
As we explained in Section 4, the decoding algorithm from [28] decodes any number of errors lower than $\operatorname{wt}(\operatorname{RM}_{d}(m))/2$. In that setting, Algorithm 1 always finishes in the first part and only requires using $D^{\mathbb{A}}_{d}(m)$. In this sense, Algorithm 1 generalizes the algorithm from [28] not only because it can correct more errors, but also because their complexity is the same when both algorithms can be applied.